# Interpolative separable density fitting on adaptive real space grids

### v1

@@ I Introduction
Even though at first sight the problem of computing and storing $V_{ijkl}$ seems too daunting for large problems, the task is made more manageable when we realize that the tensor is expected to be significantly rank deficient. Here, we show that the rank of $V_{ijkl}$, interpreted as an $N^{2}\times N^{2}$ matrix, scales linearly with the basis size $N$. As a result, the direct evaluation of (1) is typically replaced by more efficient formulations that lead to compact representations with reduced computational and storage requirements. Common approaches include resolution of identity (RI) [1, 2, 3, 4, 5], Cholesky decomposition [6, 7, 2], tensor hypercontraction (THC) [8, 9, 10, 11, 12], and the canonical polyadic (CP) approximation [13, 14, 15, 16, 17]. Other approaches include hierarchical matrix representations of the ERI [18, 19], and tensor network methods to represent the basis functions [20]. Each of these schemes involves a particular compression of the ERI tensor, with different trade-offs in terms of computational cost, accuracy, and applicability.
The most common approach used to compress the ERI tensor in quantum many-body methods, particularly in the quantum chemistry community, is the RI decomposition. It has a good balance of efficiency, accuracy and robustness, which has led to widespread use and quick adoption. It, nonetheless, suffers from serious limitations since it requires $\mathcal{O}(N^{3})$ storage and leads to $\mathcal{O}(N^{4})$ scaling algorithms when used in correlated many-body calculations. In order to reduce memory requirements and computational costs, more aggressive factorizations are needed. Among competing alternatives, THC stands out for enabling compact ($\mathcal{O}(N^{2})$ storage), low-scaling, black-box implementations of a broad range of electronic structure methods, including hybrid density functional theory (DFT) [21, 22, 23, 24], Hartree-Fock (HF) [25], coupled-cluster theories [26, 27, 28, 29, 30], many-body perturbation theory (e.g., MP2, MP3, GWGW[8, 9, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40], and auxiliary-field quantum Monte Carlo (AFQMC) [41]. One of the leading methods of constructing the THC representation is the interpolative separable density fitting (ISDF) algorithm [10, 11], which scales as $\mathcal{O}(N^{3})$, offers controllable accuracy, and relies on standard numerical linear algebra routines. Other methods of constructing the THC representation, such as least-squares (LS) THC [9, 31, 32], have $\mathcal{O}(N^{4})$ computational complexity and often require an initial RI/Cholesky decomposition. Notice that for large systems, particularly in periodic calculations with a large number of k-points, the generation of the RI/Cholesky decomposition is typically unfeasible, limiting the utility of the LS-THC approach.
@@ II THC decomposition via ISDF using adaptive grids
Here $G(\tau)$ is the imaginary time Green’s function, and $W(\tau)$ is the screened interaction. We refer to Ref. 45 for further details on the application of THC to both HF and $GW$.
@@ II.1 Interpolative separable density fitting
If one indeed discretizes $\rho_{ij}(\mathbf{r})$ on a grid of $M$ points to obtain an $M\times N^{2}$ matrix, then (7) constitutes an interpolative decomposition (ID) of that matrix [42, 43]. Several algorithms exist to construct this ID [10, 22, 46, 12]. In this work we adopt the approach described in Refs. 46, 12, summarized in App. A, which employs a pivoted Cholesky decomposition of the square of the pair density matrix. This yields the decomposition (7), with the rank $R$ chosen to meet a desired approximation accuracy.
@@
The accuracy of this decomposition is controlled by the accuracy of the ID approximation of the pair densities, which can be systematically improved by increasing $R$. Indeed, if $R=N^{2}$, then the ID can be made exact, yielding a trivial THC decomposition for which the projected Coulomb matrix contains $R^{2}=N^{4}$ degrees of freedom.
@@ II.2 Adaptive real space grids and the solution of the Poisson equation
For periodic systems, the single-particle basis functions and therefore the auxiliary basis functions are considered to be periodic, and the Poisson equation can be solved on a single unit cell with periodic boundary conditions under the charge neutrality assumption, i.e., the mean of the electron density is zero. The inner products can then be computed by the trapezoid rule with spectral accuracy [47]. For molecular systems, we assume that all single-particle basis functions have decayed sufficiently by the boundary of the simulation domain, and we solve the Poisson equation in free space. Straightforward use of the FFT to solve the Poisson equation in this setting is inappropriate, since it assumes periodic boundary conditions; in other words, it will introduce slowly-decaying spurious periodic images of the desired solution. Rather, to use uniform grids in a non-periodic setting, the FFT should be used in conjunction with a truncated kernel method, such as that of Ref. [48], which correctly imposes free space boundary conditions. In the present work, we indeed consider the molecular case, but use adaptive rather than uniform grids, so this discussion is not relevant.
@@
Many such algorithms are variants of the fast multipole method (FMM) [49, 50, 51, 52, 53, 54, 55, 56, 57] or multigrid methods [58, 59, 60, 61] (see also [62] for a comparison of the FFT, FMM, and multigrid methods). For free space problems, multigrid-based methods typically require truncating the computational domain and imposing artificial boundary conditions, so we do not consider this approach. Here, we use the recently developed dual-space multilevel kernel-splitting (DMK) framework [44]. DMK shares several characteristics with the FMM, namely its tree-based algorithmic structure and $\mathcal{O}(M)$ computational complexity for evaluating the convolution of a kernel and a function represented on an adaptive grid. However, DMK introduces significant improvements, including a simplified algorithmic framework, applicability to a broader class of kernels, and an acceleration of computationally intensive near-field calculations.
In the present work, we use adaptive octrees with product Chebyshev grids to discretize the basis functions, and the DMK algorithm, which is compatible with such a discretization, to solve the Poisson equation. Once the adaptive octree discretization has been constructed, the DMK solver can be treated as a black box, so we refer to Ref. 44 for further details on DMK.
To build the adaptive octree discretization, we begin with the single-particle basis functions $\phi_{i}(\mathbf{r})$, which are typically given in closed form. The tree is built through a recursive subdivision procedure. One begins at the root level of the tree, with a single box covering the full domain. The $\phi_{i}$ are evaluated on an $n\times n\times n$ product Chebyshev grid in the box, from which one can obtain Chebyshev polynomial interpolants $p^{(\phi_{i},B)}(\mathbf{r})=\sum_{j,k,l=0}^{n-1}C_{jkl}T_{j}(\tilde{x}_{1})T_{k}(\tilde{x}_{2})T_{l}(\tilde{x}_{3})$. Here $T_{j}$ is the degree $j$ Chebyshev polynomial of the first kind, $\mathbf{r}=(x_{1},x_{2},x_{3})$, and $\tilde{x}_{i}=2(x_{i}-c_{i})/L$ with $c_{i}$ the center of the box $B$ and $L$ its side length. The resulting approximation is accurate to order $n+1$, with a fixed, moderate value of $n$ typically chosen. If the maximum interpolation error over all $\phi_{i}$ is larger than a user-specified tolerance $\varepsilon$, the box is subdivided into eight equal boxes, and the procedure is repeated for each of the eight new boxes. Here, we use the interpolation error on box $B$ given by
@@ II.3 Summary of the algorithm and computational complexity
    Build an adaptive real space grid for the single-particle basis functions $\phi_{i}(\mathbf{r})$, with error tolerance $\varepsilon$, using the algorithm described in Sec. II.2. After upsampling, this discretization is used to represent the pair densities $\rho_{ij}(\mathbf{r})$ and the auxiliary basis functions $\zeta_{\mu}(\mathbf{r})$. The cost of this procedure scales as $\mathcal{O}(MN)$, since $N$ single-particle basis functions are discretized on a grid of $M$ points.
@@
    Using the ID procedure described in App. A, select the interpolating points $\mathbf{r}_{\mu}$ and compute the ISDF auxiliary basis functions $\zeta_{\mu}(\mathbf{r})$ on the $M$ discretization points. The cost of this procedure scales as $\mathcal{O}(R^{2}M)$.
@@
    Solve the Poisson equation for each $\zeta_{\mu}(\mathbf{r})$ on the real space grid using the DMK algorithm. The cost of this step is $\mathcal{O}(M)$ for each auxiliary basis function, or $\mathcal{O}(RM)$ in total. Compute the inner products of the results to obtain $V_{\mu\nu}$, at an $\mathcal{O}(M)$ cost per pair $(\mu,\nu)$, or $\mathcal{O}(R^{2}M)$ in total. Using (9a) and (9b), we can now form the THC decomposition (3) of the ERI tensor $V_{ijkl}$.
@@ III.2 Accuracy of ISDF truncation
In Fig. 3, for the same system as above, we plot the maximum error in $V_{ijkl}$ as a function of $\alpha$ for different values of the adaptive grid tolerance $\varepsilon$. We observe exponential convergence of the ISDF error with respect to $\alpha$ (equivalently, $R$, since $N$ is fixed), plateauing at a value determined by $\varepsilon$. Reference values of $V_{ijkl}$ are computed analytically with PySCF, which is possible due to the use of GTOs.
@@ III.3 Increasing basis set size and locality
To achieve chemical accuracy across all ERI elements, we find that $\alpha\approx 16$ is sufficient for all three basis sets. We note this value is larger than typical values used in periodic systems with pseudopotentials treated with uniform grids, for which $\alpha\approx 8$ often suffices [12]. For certain systems, we empirically observe a weak dependence of $\alpha$ on the required grid resolution, but we have not systematically studied this scaling. We also note that demanding chemical accuracy in each element of $V_{ijkl}$ is typically an unnecessary stringent criterion, since one is usually interested in observables such as the total energy and orbital energies. Furthermore, in Sec. III.5, we describe a hybrid method for electronic structure calculations which allows for smaller $\alpha$ by treating the Hartree potential term without THC.
@@ III.4 Comparison with uniform grid approach, and timings
$\varepsilon$ Setting Adaptive octree DOF Uniform grid DOF
@@
all-electron ($\sigma_{\max}\approx 3\times 10^{6}$) $715392$ Between $16384^{3}\approx 4.4\times 10^{12}$ and $32768^{3}\approx 3.5\times 10^{13}$
@@
Table 1: Number of adaptive real space grid points and estimated number of uniform grid points required to resolve $\mathrm{TiO}$ all-electron and effective core potential (ECP) basis sets for different target error tolerances $\varepsilon$.
@@
We consider two numerical setups: (i) a $\mathrm{TiO}$ molecule treated with a cc-PVTZ-ccECP basis set and an effective core potential (ECP) [68, 69], and (ii) the same system treated with an all-electron cc-PVTZ basis set [70]. For the uniform grid, the number of grid points per dimension required to achieve a target accuracy $\varepsilon$ is estimated by resolving each basis function along a representative one-dimensional slice, since resolving the localized basis functions with a uniform three-dimensional grid is intractable. In practice, we estimate the number of grid points required per dimension by doubling the grid size until the error falls below the desired tolerance, so we report the corresponding lower and upper bound estimates for each target tolerance.
@@
Table 1 compares the required grid sizes for target relative errors $\varepsilon=10^{-3}$ and $10^{-5}$. As expected, the ratio between uniform and adaptive grid sizes is several orders of magnitude larger in the all-electron setting compared to the ECP case, since the largest Gaussian exponent increases from approximately $10^{2}$ to $10^{6}$. This is because the uniform grid size increases in proportion to the largest Gaussian exponent, whereas the adaptive grid size increases only logarithmically. Still, even in the ECP case, the adaptive grid is significantly more compact than the uniform grid.
Operation Timing (sec)
Octree construction 2.67
ISDF generation of $\mathrm{r}_{\mu}$ and $\zeta_{\mu}(\mathbf{r})$ 1313.57
Poisson solves for $\zeta_{\mu}(\mathbf{r})$ 1407.59
@@
Table 2: Wall clock timings for major steps of our procedure for TiO with the all-electron cc-PVTZ basis set ($N=98$). The octree tolerance is set to $\varepsilon=10^{-3}$, and the ISDF auxiliary basis size ($R=1649$) is chosen so that the maximum error in the $V_{ijkl}$ tensor is below $10^{-3}$.
@@
Table 2 summarizes the single-core timings of the main steps of our workflow for TiO with the all-electron cc-PVTZ basis set and $\varepsilon=10^{-3}$. The cost of the octree grid construction for the $\phi_{i}$ and $\rho_{ij}$ scales only quadratically with the system size and is negligible. For this small example system, the cost of the adaptive Poisson solver step (less than one second per solve) is comparable to that of the ISDF step. We note, however, that the Poisson solver step also scales only quadratically with system size (and is trivially parallelizable over the auxiliary basis function index), whereas the ISDF step scales cubically. Thus, for larger systems, the cost of the adaptive Poisson solver will become negligible. We therefore conclude that for large systems, the cost per grid point of our fully adaptive ISDF framework is similar to that of the standard FFT-based ISDF using uniform grids, as long as the compression factor $\alpha$ does not grow significantly for systems requiring grid adaptivity (as in, for example, the experiment of Fig. 5). We note that the timing of the inner product in (9b) is not shown, since it is small compared to that of the ISDF step and has an identical cost scaling.

### v2

@@ I Introduction
Even though at first sight the problem of computing and storing $V_{ijkl}$ seems too daunting for large problems, the task is made more manageable when we realize that the tensor is expected to be significantly rank deficient. Here, we show that the rank of $V_{ijkl}$, interpreted as an $N^{2}\times N^{2}$ matrix, scales linearly with the basis size $N$. As a result, the direct evaluation of (1) is typically replaced by more efficient formulations that lead to compact representations with reduced computational and storage requirements. Common approaches include resolution of identity (RI) [1, 2, 3, 4, 5], Cholesky decomposition [6, 7, 2], tensor hypercontraction (THC) [8, 9, 10, 11, 12], and the canonical polyadic (CP) approximation [13, 14, 15, 16, 17]. Other approaches include hierarchical matrix representations of the ERI [18, 19], sparse representations [20], and tensor network methods to represent the basis functions [21]. Each of these schemes involves a particular compression of the ERI tensor, with different trade-offs in terms of computational cost, accuracy, and applicability.
The most common approach used to compress the ERI tensor in quantum many-body methods, particularly in the quantum chemistry community, is the RI decomposition. It has a good balance of efficiency, accuracy and robustness, which has led to widespread use and quick adoption. It, nonetheless, suffers from serious limitations since it requires $\mathcal{O}(N^{3})$ storage and leads to $\mathcal{O}(N^{4})$ scaling algorithms when used in correlated many-body calculations. In order to reduce memory requirements and computational costs, more aggressive factorizations are needed. Among competing alternatives, THC stands out for enabling compact ($\mathcal{O}(N^{2})$ storage), low-scaling, black-box implementations of a broad range of electronic structure methods, including hybrid density functional theory (DFT) [22, 23, 24, 25], Hartree-Fock (HF) [26], coupled-cluster theories [27, 28, 29, 30, 31], many-body perturbation theory (e.g., MP2, MP3, GWGW[8, 9, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41], and auxiliary-field quantum Monte Carlo (AFQMC) [42]. One of the leading methods of constructing the THC representation is the interpolative separable density fitting (ISDF) algorithm [10, 11], which scales as $\mathcal{O}(N^{3})$, offers controllable accuracy, and relies on standard numerical linear algebra routines. Other methods of constructing the THC representation, such as least-squares (LS) THC [9, 32, 33], have $\mathcal{O}(N^{4})$ computational complexity and often require an initial RI/Cholesky decomposition. For large systems, particularly in periodic calculations with a large number of k-points, the generation of the RI/Cholesky decomposition is typically unfeasible, limiting the utility of the LS-THC approach.
@@ II THC decomposition via ISDF using adaptive grids
Here $G(\tau)$ is the imaginary time Green’s function, and $W(\tau)$ is the screened interaction. We refer to Ref. 46 for further details on the application of THC to both HF and $GW$.
@@ II.1 Interpolative separable density fitting
If one indeed discretizes $\rho_{ij}(\mathbf{r})$ on a grid of $M$ points to obtain an $M\times N^{2}$ matrix, then (7) constitutes an interpolative decomposition (ID) of that matrix [43, 44]. Several algorithms exist to construct this ID [10, 23, 47, 12]. In this work we adopt the approach described in Refs. 47, 12, summarized in App. A, which employs a pivoted Cholesky decomposition of the square of the pair density matrix. This yields the decomposition (7), with the rank $R$ chosen to meet a desired approximation accuracy.
@@
The accuracy of this decomposition is controlled by the accuracy of the ID approximation of the pair densities, which can be systematically improved by increasing $R$. Indeed, if $R=N^{2}$, then the ID can be made exact, yielding a trivial THC decomposition for which the projected Coulomb matrix contains $R^{2}=N^{4}$ degrees of freedom. The memory requirement of the THC decomposition is dominated by that of the projected Coulomb matrix $V_{\mu\nu}$, and therefore scales as $\mathcal{O}(R^{2})$. Thus we obtain a quadratic scaling storage requirement, an improvement over the quartic scaling of the original ERI tensor, and the cubic scaling of alternative methods like RI and Cholesky decomposition.
@@ II.2 Adaptive real space grids and the solution of the Poisson equation
For periodic systems, the single-particle basis functions and therefore the auxiliary basis functions are considered to be periodic, and the Poisson equation can be solved on a single unit cell with periodic boundary conditions under the charge neutrality assumption, i.e., the mean of the electron density is zero. The inner products can then be computed by the trapezoid rule with spectral accuracy [48]. For molecular systems, we assume that all single-particle basis functions have decayed sufficiently by the boundary of the simulation domain, and we solve the Poisson equation in free space. The boundary is determined by gradually enlarging the computational domain, with an a posteriori check based on the maximum normalized $L^{2}$ norm of the single-particle basis functions evaluated on a spherical boundary. Straightforward use of the FFT to solve the Poisson equation in this setting is inappropriate, since it assumes periodic boundary conditions; in other words, it will introduce slowly-decaying spurious periodic images of the desired solution. Rather, to use uniform grids in a non-periodic setting, the FFT should be used in conjunction with a truncated kernel method, such as that of Ref. [49], which correctly imposes free space boundary conditions. In the present work, we indeed consider the molecular case, but use adaptive rather than uniform grids, so this discussion is not relevant.
@@
Many such algorithms are variants of the fast multipole method (FMM) [50, 51, 52, 53, 54, 55, 56, 57, 58] or multigrid methods [59, 60, 61, 62] (see also [63] for a comparison of the FFT, FMM, and multigrid methods, and [20] for an example of a multigrid-based approach to compute the ERI tensor). For free space problems, multigrid-based methods typically require truncating the computational domain and imposing artificial boundary conditions, so we do not consider this approach. Here, we use the recently developed dual-space multilevel kernel-splitting (DMK) framework [45]. DMK shares several characteristics with the FMM, namely its tree-based algorithmic structure and $\mathcal{O}(M)$ computational complexity for evaluating the convolution of a kernel and a function represented on an adaptive grid. However, DMK introduces significant improvements, including a simplified algorithmic framework, applicability to a broader class of kernels, and an acceleration of computationally intensive near-field calculations.
In the present work, we use adaptive octrees with product Chebyshev grids to discretize the basis functions, and the DMK algorithm, which is compatible with such a discretization, to solve the Poisson equation. Once the adaptive octree discretization has been constructed, the DMK solver can be treated as a black box, so we refer to Ref. 45 for further details on DMK.
To build the adaptive octree discretization, we begin with the single-particle basis functions $\phi_{i}(\mathbf{r})$. The tree is built through a recursive subdivision procedure. One begins at the root level of the tree, with a single box covering the full domain. The $\phi_{i}$ are evaluated on an $n\times n\times n$ product Chebyshev grid in the box, from which one can obtain Chebyshev polynomial interpolants $p^{(\phi_{i},B)}(\mathbf{r})=\sum_{j,k,l=0}^{n-1}C_{jkl}T_{j}(\tilde{x}_{1})T_{k}(\tilde{x}_{2})T_{l}(\tilde{x}_{3})$. Here $T_{j}$ is the degree $j$ Chebyshev polynomial of the first kind, $\mathbf{r}=(x_{1},x_{2},x_{3})$, and $\tilde{x}_{i}=2(x_{i}-c_{i})/L$ with $c_{i}$ the center of the box $B$ and $L$ its side length. The resulting approximation is accurate to order $n+1$, with a fixed, moderate value of $n$ typically chosen. If the maximum interpolation error over all $\phi_{i}$ is larger than a user-specified tolerance $\varepsilon$, the box is subdivided into eight equal boxes, and the procedure is repeated for each of the eight new boxes. Here, we use the interpolation error on box $B$ given by
@@
We emphasize that our adaptive grid construction procedure is agnostic to the single-particle basis set used, requiring only a basis function evaluator at arbitrary points. The adaptive grid then serves as an internal representation of the pair densities used by the DMK Poisson solver. Basis functions can be given in closed form (e.g., Gaussian-type orbitals or Slater-type orbitals), or represented in terms of values on a grid which can then be interpolated (e.g., numerical atomic orbitals [66, 67] or adaptive local basis sets [68, 69, 70, 71, 20]).
@@ II.3 Summary of the algorithm and computational complexity
    Build an adaptive real space grid for the single-particle basis functions $\phi_{i}(\mathbf{r})$, with error tolerance $\varepsilon$, using the algorithm described in Sec. II.2. After upsampling, this discretization is used to represent the pair densities $\rho_{ij}(\mathbf{r})$ and the auxiliary basis functions $\zeta_{\mu}(\mathbf{r})$. The cost of this procedure scales as $\mathcal{O}(MN)$, since $N$ single-particle basis functions are discretized on a grid of $M$ points. An $\mathcal{O}(M)$ memory cost can be obtained by building the adaptive grids for each $\phi_{i}$ sequentially by refining the grid if needed.
@@
    Using the ID procedure described in App. A, select the interpolating points $\mathbf{r}_{\mu}$ and compute the ISDF auxiliary basis functions $\zeta_{\mu}(\mathbf{r})$ on the $M$ discretization points. The cost of this procedure scales as $\mathcal{O}(R^{2}M)$, and the memory complexity as $\mathcal{O}(RM)$, corresponding to the storage of the ISDF auxiliary basis functions on the adaptive grid.
@@
    Solve the Poisson equation for each $\zeta_{\mu}(\mathbf{r})$ on the real space grid using the DMK algorithm. The cost of this step is $\mathcal{O}(M)$ for each auxiliary basis function, or $\mathcal{O}(RM)$ in total. Compute the inner products of the results to obtain $V_{\mu\nu}$, at an $\mathcal{O}(M)$ cost per pair $(\mu,\nu)$, or $\mathcal{O}(R^{2}M)$ in total. Using (9a) and (9b), we can now form the THC decomposition (3) of the ERI tensor $V_{ijkl}$. The memory required to store the matrices $X_{i\mu}$ and $V_{\mu\nu}$ in the decomposition scales as $\mathcal{O}(R^{2})$.
@@ III.2 Accuracy of ISDF truncation
In Fig. 3, for the same system as above, we plot the maximum error in $V_{ijkl}$ as a function of $\alpha$ for different values of the adaptive grid tolerance $\varepsilon$. Reference values of $V_{ijkl}$ are computed analytically using PySCF. We observe exponential convergence of the ISDF error with respect to $\alpha$ (equivalently, $R$, since $N$ is fixed), plateauing at a value determined by $\varepsilon$. This indicates that $\varepsilon$ controls the accuracy of the THC-ERI approximation, and we recommend selecting it as such in practice.
@@ III.3 Increasing basis set size and locality
To achieve chemical accuracy across all ERI elements, we find that $\alpha\approx 16$ is sufficient for all three basis sets. This value is larger than typical values used in periodic systems with pseudopotentials treated with uniform grids, for which $\alpha\approx 8$ often suffices [12]. For certain systems, we empirically observe a weak dependence of $\alpha$ on the required grid resolution, but we have not systematically studied this scaling. We emphasize that the larger values of $\alpha$ are not directly a consequence of the use of adaptive grids, but rather the locality of all-electron basis functions compared to those used in pseudopotential calculations, for which Gaussians exponents are typically approximately $10$ at most. Furthermore, demanding chemical accuracy in each element of $V_{ijkl}$ is typically an unnecessarily stringent criterion, since one is usually interested in observables such as the total energy and orbital energies. We lastly note that these larger values of $\alpha$ are required primarily to resolve sharp features associated with core electrons, which mostly contribute to the Hartree potential. In Sec. III.5, we describe a hybrid method for electronic structure calculations which allows for smaller $\alpha$ by observing that the Hartree potential term can be handled efficiently without the use of THC-ERIs.
@@ III.4 Comparison with uniform grid approach, and timings
System $\varepsilon$ Setting Adaptive octree DOF Uniform grid DOF
$10^{-3}$ ccECP ($\sigma_{\max}\approx 42$) $775872$ Between $128^{3}\approx 2.1\times 10^{6}$ and $256^{3}\approx 1.7\times 10^{7}$
$(\mathrm{NH}_{3})_{2}$ all-electron ($\sigma_{\max}\approx 1\times 10^{4}$) $854496$ Between $256^{3}\approx 1.7\times 10^{7}$ and $512^{3}\approx 1.3\times 10^{8}$
$10^{-5}$ ccECP ($\sigma_{\max}\approx 42$) $3577932$ Between $256^{3}\approx 1.7\times 10^{7}$ and $512^{3}\approx 1.3\times 10^{8}$
all-electron ($\sigma_{\max}\approx 1\times 10^{4}$) $4353588$ Between $512^{3}\approx 1.3\times 10^{8}$ and $1024^{3}\approx 1.1\times 10^{9}$
@@
$\mathrm{TiO}$ all-electron ($\sigma_{\max}\approx 3\times 10^{6}$) $715392$ Between $16384^{3}\approx 4.4\times 10^{12}$ and $32768^{3}\approx 3.5\times 10^{13}$
@@
Table 1: Number of adaptive real space grid points and estimated number of uniform grid points required to resolve effective core potential (ECP) and all-electron basis sets for $(\mathrm{NH}_{3})_{2}$ and $\mathrm{TiO}$ at different target error tolerances $\varepsilon$.
@@
For the analysis of grid sizes, we consider TiO and $(\mathrm{NH}_{3})_{2}$ molecules in two numerical setups: (i) a cc-PVTZ-ccECP basis set and an effective core potential (ECP) [75, 76], and (ii) an all-electron cc-PVTZ basis set [77]. For the uniform grid, the number of grid points per dimension required to achieve a target accuracy $\varepsilon$ is estimated by resolving each basis function along a representative one-dimensional slice, since resolving the localized basis functions with a uniform three-dimensional grid is intractable. In practice, we estimate the number of grid points required per dimension by doubling the grid size until the error falls below the desired tolerance, so we report the corresponding lower and upper bound estimates for each target tolerance.
@@
Table 1 compares the required grid sizes for target relative errors $\varepsilon=10^{-3}$ and $10^{-5}$. As expected, the ratio between uniform and adaptive grid sizes is several orders of magnitude larger in the all-electron setting compared to the ECP case, especially for $\mathrm{TiO}$, since the largest Gaussian exponent increases from approximately $10^{2}$ to $10^{6}$. This is because the uniform grid size increases in proportion to the largest Gaussian exponent, whereas the adaptive grid size increases only logarithmically. Still, even in the ECP case, the adaptive grid is significantly more compact than the uniform grid.
CnH2n+2 $n=1$ $n=2$ $n=3$ $n=4$
Octree construction 0.09 0.21 0.38 0.55
ISDF generation of $\mathrm{r}_{\mu}$ and $\zeta_{\mu}(\mathbf{r})$ 4.86 16.05 39.97 80.18
Poisson solves for $\zeta_{\mu}(\mathbf{r})$ 7.77 18.69 31.35 53.12
@@
Table 2: Wall clock timings in minutes for major steps of the adaptive ISDF procedure applied to CnH2n+2 with the all-electron aug-cc-PVDZ basis set. The octree tolerance is set to $\varepsilon=10^{-5}$, and the ISDF auxiliary basis size is fixed to $\alpha=12$. All calculations are performed using 8 OpenMP threads.
Table 2 summarizes timings of the main steps of our workflow for the alkanes CnH2n+2 using the all-electron aug-cc-PVDZ basis set and an octree tolerance $\varepsilon=10^{-5}$. The ISDF auxiliary basis size is fixed to $\alpha=12$, and all calculations are performed using 8 OpenMP threads.
@@
The cost of the octree grid construction for the $\phi_{i}$ and $\rho_{ij}$ scales only quadratically with the system size and is negligible. As the system size increases, the cost of the adaptive Poisson solver step (less than one second per solve) becomes smaller than that of the ISDF step. Indeed, the Poisson solver step also scales only quadratically with system size and is trivially parallelizable over the auxiliary basis function index, whereas the ISDF step scales cubically. Thus, for systems with a large number of orbitals, the cost of the adaptive Poisson solver will become negligible relative to that of the ISDF step.
Since the cost of the dominant ISDF step depends only on the number of grid points, and not on their spatial arrangement, we expect that for sufficiently large systems the cost per grid point of our fully adaptive ISDF framework will be similar to that of the standard FFT-based uniform grid approach, i.e., the overhead associated with grid adaptivity will become negligible. A caveat is that the required compression factor $\alpha$ might be larger for systems requiring more grid adaptivity, but our numerical experiments suggest at most mild growth (see Figs. 4 and 5). We note that the timing of the inner product in (9b) is not shown, since it is small compared to that of the ISDF step and has an identical cost scaling.